OpenStudy (stamp):

Differential Equations SPICY MATH LOTS FUN Find the general solution. y' = csc x - y cot x

OpenStudy (poison_princess1):

Here https://photomath.net/s/jYq1P

OpenStudy (stamp):

integrating factor µ = sin x ??

OpenStudy (stamp):

yeah µ = sin x and \[y (sinx)=\int cscx(sinx)=\int dx=x+C\]\[y(sinx)=x+C\]